OpenStudy (anonymous):

y varies inversely as twice x. When x = 2, y = 5. Find y when x = 10.

OpenStudy (anonymous):

y = 1

OpenStudy (anonymous):

can you help me with this one: y varies jointly with p and q. When p = 7 and q = 2, y = 70. Find y when p = 8 and q = 4.

OpenStudy (anonymous):

160

hero (hero):

Ishaan, you should show your work

OpenStudy (anonymous):

I know, Sorry y = k*p*q 70= k*2*7 k = 5 y = 5*p*q y=5*8*4 y = 160